OpenStudy (anonymous):

Use the Pythagorean Theorem to determine if the three numbers could be the measures of the sides of a right triangle. Assume that the largest is the hypotenuse. v72,17,19 *v stands for sqrt A. yes B. no

OpenStudy (anonymous):

i say no . Correct ?

OpenStudy (anonymous):

why do you say no?

OpenStudy (anonymous):

a^2 + b^2 = c^2 a = 72 b = 17

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 (sqrt(72))^2 + 17^2 = 19^2 72 + 289 = 361 361 = 361 Since the last equation is equal, the values a = sqrt(72), b = 17 and c = 19 satisfy the equation. So this is indeed a right triangle. So the answer is "yes"

OpenStudy (anonymous):

It's YES!

OpenStudy (anonymous):

That's right. 72^2 \neq 289+361\]